1.Given below is the dissociation of NH3 in water having a Kb value of 1.8 × 10-5. The reaction started with the equilibration of 0.65 M NH3. A common ion that is 0.25 M NH4Cl, was added to the initial solution and waited for some time for the reaction to reach equilibrium once again.

NH3(aq) + H2O(l) ↔ NH4+(aq) + OH-(aq)

Determine the pH of the solution given after the addition of the common ion. The Kw of water is 1.0 × 10-14.


2.What is the pH of a buffer solution that is 0.350 M ethylamine and 0.200 Methylammonium chloride, after the addition of 0.100 M HNO3? The value of Kb for ethylamine is 4.7 x 10-4.

1
Expert's answer
2022-03-08T08:03:06-0500

2. pOH = pKb + log10(acid/base)

 

pOH = 3.32 + log10(0.35/0.2)

Addition of HNO3:

pOH = 3.32 + log10(0.35+0.1/0.2)

pOH = 3.38

pH = 14 – 3.38 = 10.62